OpenStudy (anonymous):

solve by completig the square: x^2+8x=-3

OpenStudy (anonymous):

i got -4-sqrt13 and -4+sqrt13

OpenStudy (anonymous):

\[x^2 + 8x = 3 \] \[x^2 + 8x + 16 = 3 + 16 \](x + 4)62 = 19 \[x + 4 = ± \sqrt{19}\] \[x = -4 ± {\sqrt{19}}\]

OpenStudy (anonymous):

thanks